OpenStudy (anonymous):

An internet bookseller charges $6 perpaperback book plus a $3 shipping and handling fee per order.if Mr.Montoya spent $111 on books for his class,how many books did he buy?

OpenStudy (anonymous):

this is a bit of hunt and peck method. each book cost $6 plus the $3 shipping fee. Keeping this in mind gives a consistent breakdown of expenses. 6 x #of books + 3 x #of books for shipping = 111 so you would substitute and find the answer 6 x 10 + 3 x 10 = NOT ENOUGH 6 x 15 + 3 x 15 = TOO MUCH 6 x 11-14 + 3 x 11-14 = gives the answer! Hope this helps you out

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

if we let the number of books bought by Mr. Montoya be x, then expressing the given information in an equation 6x + 3 = 111 (i.e 6x being the price he would have to pay for all the books he bought and 3 being the shipping fee assuming he placed one order for all of the books) solving the obtained equation, 6x + 3 =111, =6x + 3 - 3 = 111 - 3 =6x = 108, divide both sides of the equation by 6, then x = 18 therefore Mr. Montoya bought 18 books

OpenStudy (anonymous):

fatma don't forget that it is $3 per order ... 18*3=54 this would undo your solution to the problem

OpenStudy (anonymous):

in answering the question i assumed that he placed one order for all of the books plus it is as you said $3 per order, not per book